Show that if an integer n is not divisible by 3, then n2 – 1 must be a multiple of 3


1
Expert's answer
2022-05-06T05:53:23-0400

If "n" is not divisible by 3, then either "n=3k+1, k\in \Z" or "n=3m+2," "m\in \Z."

Then


"n^2-1=(3k+1)^2-1=9k^2+6k"

"=3(3k^2+2k), k\in \Z"

Let "p=3k^2+2k." We know that "p" is an integer because "k" is an integer. Therefore "n^2-1" equals 3 times some integer. So "n^2-1" is a multiple of 3.

Or


"n^2-1=(3m+2)^2-1=9m^2+12m+3"

"=3(3m^2+4m+1), m\in \Z"

Let "q=3m^2+4m+1." We know that "q" is an integer because "m" is an integer. Therefore "n^2-1" equals 3 times some integer. So "n^2-1" is a multiple of 3.

Therefore if an integer "n" is not divisible by 3, then "n^2 \u2013 1" must be a multiple of 3.